a collection of nickels dimes and quarters in a money bag worth 33.50 number of quarters is five times the number of nickels and the number of dimes is five less than four times the number of nickels find the number of each type of coin.
Answer by ptaylor(1894)

Let x=the number of nickels; Value of the nickels=0.05x
Then 5x=number of quarters; Value of the quarters=0.25(5x) or 1.25x
And 4x-5=number of dimes; Value of the dimes=0.10(4x-5)
Now we are told that the value of these coins adds up to $33.50 so our equation to solve is:
0.05x+1.25x+0.10(4x-5)=$33.50 get rid of parens
0.05x+1.25x+0.4x-0.5=33.50 add 0.5 to both sides
0.05x+1.25x+0.4x-0.5+0.5=33.50+0.5 collect like terms
1.7x=34 divide both sides by 1.7
x=20----------------number of nickels
5x=5*20=100-----------------number of quarters
4x-5=4*20-5=75-----------number of dimes
CK
0.05*20+0.25*100+0.10*75=33.50
1+25+7.50=33.50
33.50=33.50
Hope this helps---ptaylor